OpenStudy (anonymous):

With the given lengths would this be a right triangle? (square root symbols) √11 yds, √5 yds, 16yds

OpenStudy (anonymous):

pythagorean theorem

OpenStudy (anonymous):

what do you know about Pythagorean theorem?

OpenStudy (anonymous):

a squared + b squared = c squared

OpenStudy (anonymous):

then why don't you try the combination and see if they satisfy the theorem.

OpenStudy (anonymous):

and make it equal 180 degrees for a triangle?

OpenStudy (anonymous):

square root off is 3.3

OpenStudy (anonymous):

I meant square root of 11 is 3.3

OpenStudy (anonymous):

huh? why are you making the problem more complicated than it is? They have given you lengths of the three sides of a triangle. All they want to know is if the triangle is a right angled triangle. The way to find out if a triangle is right angled or not is to see if its sides satisfy the pythagorean theorem. That is it. Don't try and make it harder for yourself.

OpenStudy (anonymous):

ok so two square roots and one is not... therefore no

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[(\sqrt{11})^{2}+(\sqrt{5})^{2} = ?\]

OpenStudy (anonymous):

Ok I'm not sure... but I came up with 6.6 + 4.4 = 11 by definition: the square of the length of the hypotenuse is equal to the sum of the squares of the lengths of the legs. So again I would say this is not a right triangle ??

OpenStudy (anonymous):

yes, you are right. It is not a right angle triangle.

OpenStudy (anonymous):

was that right how I did the calculations tho? took square root of 11 times it by two .... the same for square root of 5 times by two and then add ... was that number supposed to add up to 16? to make it a right triangle?

OpenStudy (anonymous):

wait...3.3 squared plus 2.2 squared = 10.89 + 4.84 = 15.73 thats close to 16

OpenStudy (anonymous):

square root of 4 is 2. that means 2 times 2 is 4. that means \[\sqrt{4} \times \sqrt{4} = 4\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so \[(\sqrt{11})^{2} = \sqrt{11} \times \sqrt{11} = 11\]

OpenStudy (anonymous):

oh I surely did not know that...

OpenStudy (anonymous):

the very fact that you are squaring square roots of a number means that you will get the number within the square root. That is the definition of square root.

OpenStudy (anonymous):

what did you think square root meant then?

OpenStudy (anonymous):

I don't know... I do know the square root of 81 is 9...

OpenStudy (anonymous):

stuff like that

OpenStudy (anonymous):

yes, square root of 81 is 9. that means that 9 times 9 is 81. similarly square root of 3 is 1.73205081 that means that 1.73205081 times 1.73205081 is 3.

OpenStudy (anonymous):

instead of calling it square root everytime, we write it is \[\sqrt{3}\]

OpenStudy (anonymous):

yes I know... not sure how to do symbols on line

OpenStudy (anonymous):

11 + 5 = 16 therefore it is a right triangle

OpenStudy (anonymous):

what is the third side? is it 16 or square root(16)

OpenStudy (anonymous):

just 16

OpenStudy (anonymous):

square root of 16 I know is 4

OpenStudy (anonymous):

so what does the pythagorean theorem say?

OpenStudy (anonymous):

I'm confused...... I already said it... I just dont know if these numbers are supposed to add up to something or have all perfect numbers from the roots? UGGGH

OpenStudy (anonymous):

or the two square roots equal 16? Please be patient with me....

OpenStudy (anonymous):

\[(\sqrt{11})^{2} + (\sqrt{5})^{2} = 16 \neq 16^{2}\]

OpenStudy (anonymous):

I'm ready to pull my hair out!!!

OpenStudy (anonymous):

Thank you. That makes sense... that was the part I wasnt sure of.

OpenStudy (anonymous):

Here is another one.... (√7)2 + (√2)2 = 9 does not equal √9 Not a right triangle

OpenStudy (anonymous):

Got it and thank you for your help!!!!!